# Thermodynamic and observational implications of black holes in toroidal geometry

### v2

@@ VI.2 Null Particles
We present the behavior of redshift $z_{1}$ (upper left panel), blueshift $z_{2}$ (upper right panel) and gravitation shift $z_{\mathrm{g}}$ (lower panel) by setting $M=1,~r_{\mathrm{e}}=r_{\mathrm{ISCO}}$ and compute various curves for different values of $Q$ such as for $Q=0.8$ (red curve), $Q=1$ (blue dotted curve) and $Q=1.3$ (black dotted-dashed curve). It is evident from Fig. 9 that raising the parameter $Q$ leads to higher to lower values of the redshift $z_{1}$ (left panel) (right panel), while in the case of blueshift $z_{2}$ (middle panel), it increases with the lower charge $Q$. It also indicates that, as the parameter $Q$ increases, the frequency-shift curves grow more rapidly with $l$. It is worthy to mention here that $z_{\mathrm{g}}$ increases with neagtive behavior after the AdS radii $l$ increases it becomes positive and rises with the parameter $Q$ as presented in Fig. 9 (lower panel) while on the other hand $z_{2}$ (middle panel) shows negative trend which becomes more negative as the parameter $Q$ rises, owing to the impact of $z_{\mathrm{kin}^{-}}$ component. Furthermore, from the relation of gravitation shift, we observe that it is independent of the term $l$, so we graphically present the behavior of $z_{\mathrm{g}}$ in the form of $r_{\mathrm{e}}$, which indicates that for large values of $r_{\mathrm{e}}$, gravitational shift obtains maximum values. Thereby, the curves in Fig. 9 represent the highest frequency shifts, computed for an emitter revolving circularly around the BH at $r_{\mathrm{e}}=r_{\mathrm{ISCO}}$, and also as the orbital radius grows infinitely, then the frequency shift vanishes to zero. One can plugged $R=1+z_{1}$ and $B=1+z_{2}$ to verify
@@ VII Conclusions
In this work, we have comprehensively discussed the thermodynamic and observational analysis for a charged torus-like BH. This paper is structured in two parts; the initial part focuses on the impact of different entropy frameworks, such as the Hawking-Bekenstein entropy, exponential corrected, and Rènyi entropies, on the thermodynamics of a charge torus-like BH, while the latter part focuses on the observational aspects determined purely from the spacetime geometry. In the thermodynamic analysis, we have employed different entropy frameworks, such as the Hawking-Bekenstein, exponential corrected, and Rènyi entropies, to analytically compute various thermodynamic quantities such as mass $M$, temperature $T$, and thermodynamic volume $V$. By utilizing these quantities, we computed and graphically presented the heat capacity, PVP-V diagram, isotherm compressibility, and free energies (HFE and GFE).
Table 2: This table presents the thermodynamic quantities by employing the Hawking-Bekenstein, Rényi, and exponential corrected entropies for the charged torus-like BH, and also its comparison with the thermodynamic quantities of the RN-Ads and Schwarzschild BHs. Here, we present the thermodynamic quantities like temperature, heat capacity, isothermal compressibility, HFE, GFE, and Ruppeiner Ricci scalar by $C,~K_{T}~\mathcal{F},~\mathcal{G}$, and $R^{\mathrm{Rup}}$, respectively.
BHs CTL BH RN-AdS Schwarzschild
Quantitites HBE Rényi ECE HBE HBE
$C$ UnStable Unstable UnStable Unstable Unstable
ZPs 1 1 2 1 0
PVP-V Monotonic Monotonic Non-monotonic Non-monotonic not defined
$K_{T}$ Positive Positive Divergence Divergence not defined
$\mathcal{F}$ UnStable Unstable Stable Unstable Unstable
$\mathcal{G}$ Stable Unstable Stable Stable Stable
$R^{\mathrm{Rup}}$ Coincide Coincide Coincide both ZPs Coincide No-Coincidence
@@
In this analysis, it is notably found that there is only one ZP of heat capacity in the case of the Hawking-Bekenstein and Rènyi, while on the other hand, the exponential corrected entropy demonstrates two distinct ZPs of heat capacity. This suggests that exponentially corrected entropy provides a more intricate thermal phase structure and more complex stability behavior. To validate the physical significance of these identified ZPs, we employed Weinhold and Ruppeiner geometries, observing that the Ricci scalar obtained from the Ruppeiner geometry diverges at the same point where heat capacity becomes zero, thereby providing geometric validation of the phase transition. Compared to the Hawking-Bekenstein and Rènyi entropies, where free energies (HFE and GFE) demonstrate the fluctuating behavior indicating instability, the exponential corrected entropy approach leads to positive and stable behavior in both HFE and GFE, highlighting global thermodynamic stability. Similarly, the exponential corrected entropy model shows a non-monotonic pressure trend in the PVP-V diagram, indicating a possible phase transition absent in other entropy models, and this is also confirmed by the behavior of isothermal compressibility. Additionally, we presented the impact of entropy corrections on the thermodynamics of charged torus-like BH by comparing it with the spherically symmetric spacetimes such as Schwarzschild and RN-AdS BHs in table 2.
@@
We further studied the sparsity of Hawking radiation and the energy emission rate by utilizing the conjugate temperature, which corresponds to the respective entropy models used in this analysis. It is observed that for small ranges of entropy, the parameter of sparsity $\eta$ increased, but for large values of entropy, it began to decrease. After some time, it became constant for the exponential corrected entropy, but in the case of the Hawking-Bekenstein and Rènyi entropy, it converges to zero. As we mentioned earlier, our aim in this study is to investigate the theoretical (thermodynamics) and observational characteristics of the charged torus-like BH. We analyzed the motion of massive and massless test particles via geodesics to examine the observational aspects of a charged torus-like BH. We analytically determined the redshift, blueshift, and gravitational frequency shift of emitted photons originating from the orbital motion. It is notably observed that, for specific parameter choices such as $Q>1$ and $M\leq 1$, the condition $r_{\mathrm{ph}}<r_{\mathrm{ISCO}}$ may not hold for charged torus-like BHs due to their spacetime geometry. We observed that this inequality might hold under the following conditions, such as $M>1$ and $Q\leq 1$. Additionally, our analysis confirms that the calculated redshift $z_{1}$ and blueshift $z_{2}$ agree with the observational constraints for the megamaser disk systems NGC 4258 and UGC 3789. The observed consistency suggests that toroidal BH models may be suitable for studying specific astrophysical BHs, notably when examined through the lens of frequency-shift phenomena. Our analysis highlights the significant role of entropy correction in BH thermodynamics and identifies observational features arising from toroidal geometries.
@@ Acknowledgements.
The work of KB was supported by the JSPS KAKENHI Grant Number 24KF0100 and Competitive Research Funds for Fukushima University Faculty (25RK011).

### v3

@@ VI.2 Null Particles
We present the behavior of redshift $z_{1}$ (upper left panel), blueshift $z_{2}$ (upper right panel) and gravitation shift $z_{\mathrm{g}}$ (lower panel) by setting $M=1,~r_{\mathrm{e}}=r_{\mathrm{ISCO}}$ and compute various curves for different values of $Q$ such as for $Q=0.8$ (red curve), $Q=1$ (blue dotted curve) and $Q=1.3$ (black dot-dashed curve). It is evident from Fig. 9 that increasing the parameter $Q$ leads to a transition from higher to lower values of the redshift $z_{1}$ (upper left panel), whereas in the case of blueshift $z_{2}$ (upper right panel), it increases for different values of charge $Q$. It also indicates that, as the parameter $Q$ increases, the frequency-shift curves grow more rapidly with $l$. Notably, the gravitational shift $z_{\mathrm{g}}$ is initially negative but turns positive and increase with $Q$ as the AdS radii $l$ grows as shown in Fig. 9 (lower panel), while on the other hand $z_{2}$ (upper right panel) becomes increasingly negative with rising $Q$, owing to the impact of $z_{\mathrm{kin}^{-}}$ component. Furthermore, from the relation of gravitation shift, we observe that it is independent of the term $l$, so we graphically present the behavior of $z_{\mathrm{g}}$ in the form of $r_{\mathrm{e}}$, which indicates that for large values of $r_{\mathrm{e}}$, gravitational shift obtains maximum values. Thereby, the curves in Fig. 9 represent the highest frequency shifts, computed for an emitter revolving circularly around the BH at $r_{\mathrm{e}}=r_{\mathrm{ISCO}}$, and also as the orbital radius grows infinitely, then the frequency shift vanishes to zero. One can plugged $R=1+z_{1}$ and $B=1+z_{2}$ to verify
@@ VII Conclusions
In this work, we have comprehensively discussed the thermodynamic and observational analysis for a charged torus-like BH. This paper is structured in two parts; the initial part focuses on the impact of different entropy frameworks, such as the Hawking-Bekenstein entropy, exponential corrected, and Rényi entropies, on the thermodynamics of a charge torus-like BH, while the latter part focuses on the observational aspects determined purely from the spacetime geometry. In the thermodynamic analysis, we have employed different entropy frameworks, such as the Hawking-Bekenstein, exponential corrected, and Rényi entropies, to analytically compute various thermodynamic quantities such as mass $M$, temperature $T$, and thermodynamic volume $V$. By utilizing these quantities, we computed and graphically presented the heat capacity, PVP-V diagram, isotherm compressibility, and free energies (HFE and GFE). In this analysis, it is notably found that there is only one ZP of heat capacity in the case of the Hawking-Bekenstein and Rényi, while on the other hand, the exponential corrected entropy demonstrates two distinct ZPs of heat capacity. This suggests that exponentially corrected entropy provides a more intricate thermal phase structure and more complex stability behavior. To validate the physical significance of these identified ZPs, we employed Weinhold and Ruppeiner geometries, observing that the Ricci scalar obtained from the Ruppeiner geometry diverges at the same point where heat capacity becomes zero, thereby providing geometric validation of the phase transition. Compared to the Hawking-Bekenstein and Rényi entropies, where free energies (HFE and GFE) demonstrate the fluctuating behavior indicating instability, the exponential corrected entropy approach leads to positive and stable behavior in both HFE and GFE, highlighting global thermodynamic stability. Similarly, the exponential corrected entropy model shows a non-monotonic pressure trend in the PVP-V diagram, indicating a possible phase transition absent in other entropy models, and this is also confirmed by the behavior of isothermal compressibility. Additionally, we presented the impact of entropy corrections on the thermodynamics of charged torus-like BH by comparing it with the spherically symmetric spacetimes such as Schwarzschild and RN-AdS BHs in Table 1.
Furthermore, analysis through the GTPE framework clarifies that exponentially corrected entropy exhibits significant thermodynamic deviations, highlighting its strong sensitivity to microstructures of the BHs as summarized in Table 1. Whereas, the Hawking-Bekenstein and Rényi entropies align well with GPTE, signalling the same universality class; however, the exponential corrected entropy behaves differently, reflecting a non-perturbative modification. This highlights that, as compared to other entropy models, the exponentially corrected entropy uniquely captures subtle microstructural aspects linked to the toroidal horizon topology. Our analysis reveals that the choice of entropy plays a pivotal role in determining the thermodynamics of BHs in toroidal spacetime, highlighting the ability of most generalized entropies to investigate quantum and topological phenomena beyond the traditional area law. Thereby, this study offers a robust and physically motivated groundwork for examining non-perturbative and non-extensive entropy corrections in BH thermodynamics, and also suggests various possible directions for future work utilizing the generalized entropy (with six, five, or four parameters) framework as given in Refs. Nojiri:2022aof ; Nojiri:2022dkr ; Nojiri:2022sfd ; Elizalde:2025iku ; Odintsov:2023vpj ; Nojiri:2024zdu ; Odintsov:2022qnn . For example, a worthwhile future direction is to employ the generalized six parameters entropy via numerical inversion methods to further examine the universality among topological BHs, which could provide more valuable insights.
@@
We further studied the sparsity of Hawking radiation and the energy emission rate by utilizing the conjugate temperature, which corresponds to the respective entropy models used in this analysis. It is observed that for small ranges of entropy, the parameter of sparsity $\eta$ increased, but for large values of entropy, it began to decrease. After some time, it became constant for the exponential corrected entropy, but in the case of the Hawking-Bekenstein and Rényi entropy, it converges to zero. As we mentioned earlier, our aim in this study is to investigate the theoretical (thermodynamics) and observational characteristics of the charged torus-like BH. We analyzed the motion of massive and massless test particles via geodesics to examine the observational aspects of a charged torus-like BH. We analytically determined the redshift, blueshift, and gravitational frequency shift of emitted photons originating from the orbital motion. It is notably observed that, for specific parameter choices such as $Q>1$ and $M\leq 1$, the condition $r_{\mathrm{ph}}<r_{\mathrm{ISCO}}$ may not hold for charged torus-like BHs due to their spacetime geometry. We observed that this inequality might hold under the following conditions, such as $M>1$ and $Q\leq 1$. Additionally, our analysis confirms that the calculated redshift $z_{1}$ and blueshift $z_{2}$ agree with the observational constraints for the megamaser disk systems NGC 4258 and UGC 3789. The observed consistency suggests that toroidal BH models may be suitable for studying specific astrophysical BHs, notably when examined through the lens of frequency-shift phenomena. Our analysis highlights the significant role of entropy correction in BH thermodynamics and identifies observational features arising from toroidal geometries.
@@ Acknowledgements.
The work of K.B. was supported by the JSPS KAKENHI Grant Numbers 24KF0100, 25KF0176, and Competitive Research Funds for Fukushima University Faculty (25RK011).